OpenStudy (a1234):

How do you find the molar mass when given a mass per atom value as 15.99 u/atom?

OpenStudy (michele_laino):

one mol contains: \[{N_A} = 6.02 \cdot {10^{23}}{\text{ atoms}}\] so we have to do this computation: \[m{\text{ = 15}}{\text{.99}} \cdot 6.02 \cdot {10^{23}}=...?\] \(N_A\) is the \(Avogadro\)'s number

OpenStudy (a1234):

I mean 96.26 X 10^23 g/mol

OpenStudy (michele_laino):

yes!

OpenStudy (a1234):

That's it?

OpenStudy (michele_laino):

yes! I think so!

OpenStudy (a1234):

Thanks!

OpenStudy (michele_laino):

:)